OpenStudy (anonymous):

Solve Whole Sqrt of ( 28 - 10 sqrt 3 )

OpenStudy (raden):

if it just simplifying form, you can use the formula : sqrt{(a+b) + 2sqrt(a*b)} = sqrt(a) + sqrt(b) with a > b so, Sqrt ( 28 - 10 sqrt(3) ) = sqrt of ( 28 - 2 * 5 sqrt(3) ) = sqrt of ( 28 - 2 * sqrt(25) * sqrt(3) ) = sqrt of ( 28 - 2 * sqrt (25 * 3 )) = sqrt of ( (25+3) - 2 * sqrt (25 * 3) = sqrt(25) + sqrt(3) = 5 + sqrt(3)